OpenStudy (theviper):

\(\LARGE{\color{red}{Tutorial:-}}\) How to find the value of 'x' & 'y' in the equations by Elimination Method........

OpenStudy (theviper):

Example:- Solve for 'x' & 'y' : 10x+3y=75 ; 6x-5y=11 Solution:- The given equations are: 10x+3y=75 ..................(i) 6x-5y=11 .................(ii) Multiplying (i) by 5 & (ii) by 3, we get: 50x+15y=375 ..................(iii) 18x-15y=33 ..................(iv) Adding (iii) & (iv), we get: 60x=408 \(\Large{\implies x=\frac{408}{68}=6}\) Putting x=6 in (i), we get \(\Large{(10\times6)+3y=75\implies 60+3y=75}\) \(\Large{\implies 3y=(75-60)}\) \(\Large{\implies 3y=15\implies y=5.}\)

OpenStudy (sasogeek):

x?

OpenStudy (theviper):

6

OpenStudy (unklerhaukus):

how did you get 60x=408 ?

OpenStudy (unklerhaukus):

50x+15y=375 ..................(iii) 18x-15y=33 ..................(iv) Adding (iii) & (iv), we get: 68x=408

OpenStudy (theviper):

Adding as u see @UnkleRhaukus :) 50x+15y=375 +18x-15y=+33 -------------------------- 68x=408

OpenStudy (unklerhaukus):

ok , it just a minor typo

OpenStudy (theviper):

:)

OpenStudy (unklerhaukus):

i'd finish a simultaneous equations question with testing the solution we found in both the equations \[(x,y)=(6,5)\mapsto10x+3y=75\] \[10(6)+3(5)=75\]\[60+15=75\]\[75=75\qquad \text {true}\] \[(x,y)=(6,5)\mapsto 6x-5y=11\] \[6(6)-5(5)=11\]\[36-25=11\]\[11=11\qquad \text{true}\]

OpenStudy (theviper):

Right

OpenStudy (theviper):

It can be used to check if our \(\Large{\cal{Answer}}\) is right or not ;)

OpenStudy (theviper):

Thanx for that @UnkleRhaukus in my Tutorial :)

OpenStudy (anonymous):

we could also just be very difficult and use crammers rule to work it out.

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

what is good

OpenStudy (theviper):

my tutorial I think :P

OpenStudy (theviper):

:)

mathslover (mathslover):

\[\huge{\color{blue}{\cal{G}}\color{red}{\textbf{R}}\color{green}{\mathsf{E}}\color{brown}{\frak{A}}\color{orange}{\mathbb{T}}\space \color{blue}{\textbf{T}}\color{black}{\cal{U}}\color{orange}{\mathbb{T}}\color{blue}{\frak{O}}\color{green}{\mathbb{R}}\color{red}{\mathsf{I}}\color{green}{\frak{A}}\color{brown}{\frak{L!!!}}}\\\]

OpenStudy (anonymous):

being on openstudy.com with all guys.. I find that a very good thing myself but hey that is just me.

OpenStudy (theviper):

Wow so nice \(LATEX\):)

mathslover (mathslover):

\[\huge{\color{blue}{\cal{:D thanks}}\space \color{green}{\frak{TheViper}}}\]

OpenStudy (theviper):

\[\Large{\color{red}{\cal{Great \space \color{green}{L}\color{blue}{A}\color{pink}{T}\color{gold}{E}\color{purple}{X}}}}\]

OpenStudy (anonymous):

Viper your da man

mathslover (mathslover):

\[\huge{\color{grey}{\frak{TH}}\color{green}{\cal{A}}\color{orange}{\mathbb{N}}\color{blue}{\mathsf{K}}\color{brown}{\textbf{S}}\space \color{violetred}{\textbf{Again TheViper}}}\]

mathslover (mathslover):

now I stop this latex making and will not disturb your tutorial thanks

OpenStudy (theviper):

o.O wow @mathslover {Your wish} But my advice is that plz give a tutorial on LaTeX :) U the Best !!!

OpenStudy (theviper):

yw

mathslover (mathslover):

there is already a tutorial for latex...

OpenStudy (anonymous):

is this the new math or something?? because I wasn't told about this. lol

OpenStudy (anonymous):

I am a dumb man in using latex :S

OpenStudy (anonymous):

May be one day Latex 5 I will learn as well. May be there is a College of Latex out there somewhere lol